An object's moment of inertia is 2.0kg?m2 . Its angular velocity is increasing at the rate of 3.8rad/s per second.

What is the net torque on the object?
Answer #1

we know,

torque= momentInertia*angularacceleration

=> torque= 2.0 x 3.8

=>torque =7.6 N-m
Answer #4

here,

angular acceleration , a = 3.8rad/s^2

then

Torque = I * a = 2 * 3.8 = 7.6 N.m